OpenStudy (anonymous):

What is the solution to the following equation? 5(2x − 6) + 20 = 10 9 5 3 2

pooja195 (pooja195):

We did one like this try solving it yourself tell me what u come up with

OpenStudy (anonymous):

4

pooja195 (pooja195):

?

OpenStudy (anonymous):

5*

pooja195 (pooja195):

Try again. Please show your work

pooja195 (pooja195):

5(2x−6)+20=10 5(2x−6)+20=10 5(2x−6)+20=10 (5)(2x)+(5)(−6)+20=10(Distribute) 10x+−30+20=10 (10x)+(−30+20)=10(Combine Like Terms) 10x+−10=10 10x−10=10 10x−10+10=10+10 10x=20 Take it from here

OpenStudy (anonymous):

I don't like math so

pooja195 (pooja195):

yea? Well i dont like when people dont try.

OpenStudy (anonymous):

the answer is 9

pooja195 (pooja195):

Whats 20/10 ?